# Maximum Independent Set; Hardness Parameter; Neutral Atoms; Rydberg Blockade; QuEra’s Aquila; Analog Quantum Computing; Adiabatic Schedules; Counterdiabatic Driving

### v2

@@ Abstract
We propose a numerical approach to design highly efficient adiabatic schedules for analog quantum computing, focusing on the maximum independent set problem and neutral atom platforms. Based on a representative dataset of small graphs, we present numerical evidences that the optimum schedules depend principally on the hardness of the problem rather than on its size. These schedules perform better than the benchmark protocols and admit a straightforward implementation in the hardware. This allows us to extrapolate the results to larger graphs and to successfully solve moderately hard problems using QuEra’s 256-qubit Aquila computer. We believe that extending our approach to hybrid algorithms could be the key to solve the hardest instances with the current technology, making yet another step toward real-world applications.
1 Introduction
@@
In recent years, quantum computers have emerged as a promising way to solve complex optimization tasks in science and industry [1, 2, 3, 4]. Typically, the problem of interest is encoded in the ground state of a quantum Hamiltonian, which can be reached through adiabatic evolution starting from an easy-to-prepare initial state [5, 6, 7, 8]. Advanced protocols have to be developed to overcome not only the experimental limitations, such as qubit decoherence and control errors, but also the intrinsic complication of vanishing gap in the eigenspectrum during the evolution [9, 10, 11, 12].
@@
Here, we focus on the maximum independent set (MIS) problem, where the task is to find the largest set of vertices that do not share an edge in a graph. Such problems appear naturally in practical applications, for instance while designing resilient communication networks [13], and are known to be NP-hard in general [14]. Finding a MIS is still a difficult task in the particular case of unit disk graphs [15, 16], in which vertices are connected by an edge if they are closer than a fixed distance. This family of graphs admits a direct mapping to arrays of Rydberg atoms, which explains the popularity of the MIS problem in analog quantum computing [17, 18, 19]. In that setting, optical tweezers enable arbitrary arrangements of atoms and controlled laser pulses manipulate the quantum states [20, 21]. The lasers then drive the evolution of the system according to schedules that are specific to the problem to be solved.
@@
The precise description of the adiabatic schedules is of utmost importance since it conditions the success of the quantum algorithm. Many methods have been proposed in this regard: quantum approximate optimization algorithm and variational quantum eigensolver [22, 23, 24], shortcut to adiabadicity and counterdiabatic protocol [25, 26, 27, 28], Bayesian optimization and machine learning [29, 30], to name a few. Most aforementioned methods require multiple iterations to be run on a quantum device and to be updated by a classical optimizer and thus belong to the class of hybrid algorithms [31]. By their very nature, these algorithms cost both time and money as they make intense use of the quantum hardware.
In this study, we show how a precomputation of optimized schedules for small graphs can lead to highly efficient protocols for graphs of larger size on the quantum device [^1] We equally refer to the number of vertices in a graph as order or size. In quantum computing, the latter is commonly understood as the size of the system and not as the number of edges in the graph (mathematical definition).. First, we introduce the basic concepts that are used throughout this work. Then, in Sec. 2 we describe how to optimize the schedules based on numerical simulations of small graphs, and in Sec. 3 we present the results obtained in experiments. Detailed information and calculations are available in the appendix.
Rydberg Hamiltonian
@@
The analog quantum dynamics is governed by the Hamiltonian $H=H_{\text{drive}}+H_{\text{cost}}$, where the driving part and the cost function are [19]
@@
$\frac{H_{\text{cost}}(t)}{\hbar}$ $=-\Delta(t)\sum_{i}\hat{n}_{i}+\sum_{i<j}V_{ij}\hat{n}_{i}\hat{n}_{j}.$ (1b)
In these equations, $\Omega$ is the Rabi drive amplitude, $\phi$ its phase, $\Delta$ the detuning, and $\hat{n}_{i}=\ket{1}_{i}\bra{1}$ | detects the Rydberg excitation of the atom $i=1,\ldots,N$. The two-body potential is $V_{ij}=C_{6}/|\textbf{r}_{i}-\textbf{r}_{j}|^{6}$, with $C_{6}$ a constant for the Van der Waals interaction between two Rydberg states at positions $\textbf{r}_{i}$ and $\textbf{r}_{j}$, encoding the unit disk graph of the MIS problem. The boundary conditions at initial and final times $t_{\text{i}}$ and $t_{\text{f}}$ for the adiabatic evolution are usually stated as follows:
@@
In practice, we implement the first inequality as $\Delta_{\text{i}}\leq-\delta_{\Delta}$, where $\delta_{\Delta}$ is a small positive constant that incorporates the main sources of error in the quantum hardware, so that all atoms are ensured to start in the ground state |$\ket{0}$. More importantly, the final detuning should lie in a narrow interval $[\Delta_{\text{LB}},\Delta_{\text{UB}}]$, where the bounds are explicitly given by
$\frac{C_{6}}{12\,a^{6}}\leq\Delta_{\text{f}}\leq\frac{C_{6}}{8\,a^{6}}$ (3)
in the case of an underlying square lattice with a distance $a$ between adjacent nodes, see Appendix A. Note that such bounds are not universal, but they are satisfied by the vast majority of unit disk graphs for a given lattice.
Quantum hardware
@@
Hardness parameter
A detailed analysis of the classical dynamics of the simulated annealing for MIS reveals that the graphs with many suboptimal solutions are likely to trap the algorithm in local minima, hence missing the global solution [19]. This observation leads to the following definition of the hardness parameter of a graph:
$\mathcal{HP}=\frac{D_{|\text{MIS}|-1}}{|\text{MIS}|\cdot D_{|\text{MIS}|}},$ (4)
[…]
2 Numerical optimization
We consider a fixed set of 500 representative graphs of size 8 to 17, which allows us to perform fast numerical simulations and to use efficient deterministic optimization algorithms, see Appendix C. This dataset is such that the correlation between the size of the graphs and their hardness parameter is minimized. Therefore, the results that depend principally on $\mathcal{HP}$ can be extrapolated, to a certain extent, to graphs of any size.
@@ 2 Numerical optimization
where |$\ket{\psi(\Theta,t_{\text{f}})}$ is the final quantum state and |$\ket{\text{MIS}_{i}}$ is the basis vector that corresponds to the $i$-th solution of the MIS problem. Note that most studies consider the approximation ratio instead (normalized mean energy of the final state), but in our opinion only the lowest energy level and not the full energy spectrum has a practical utility because of the very definition of the hardness parameter in Eq. (4).
Following the “best practices” listed in [21], we consider protocols for which the Rabi amplitude $\Omega(t)$ reaches $\Omega_{\max}$ during the evolution, so that the associated dynamic blockade radius becomes $R_{\text{b}}=(C_{6}/\Omega_{\max})^{1/6}\approx 8.4$. In order to generate a unit disk graph on the nearest and next-nearest neighbors of a square lattice, this radius should satisfy $\sqrt{2}a<R_{\text{b}}<2a$, and consequently $5.94.2<a<$ 5.9. We set $a=5$, which is close to the geometric mean of the bounds and which yields good results in preliminary calculations. We consider a fixed evolution time $t_{\text{f}}=1$ so that $P_{\text{MIS}}$ varies as much as possible depending on the various protocols; much smaller or larger times lead to either extremely bad or extremely good results independently of the precise shape of the schedules.
@@
2.1 Piecewise linear protocol
The typical adiabatic protocol is based on piecewise linear schedules, where $\Omega$ is slowly ramped up, then $\Delta$ is ramped from negative to positive, and finally $\Omega$ is ramped down, see [21, Fig. 3.1b] or [35, Fig. 2c]. Usually, these schedules are symmetric in the sense that the initial and final ramp times $\tau_{\text{i}}$ and $\tau_{\text{f}}$ are equal, and that the initial and final detunings $\Delta_{\text{i}}$ and $\Delta_{\text{f}}$ have opposite values.
In our case, we are looking for schedules that are as general as possible, so we keep all parameters independent. Moreover, we consider distinct values of the ramp times for $\Omega$ and $\Delta$. After few optimization steps, it appears that the ramp times of the detuning tend to vanish; we therefore set them to zero. Then, we optimize the four remaining parameters $(\tau_{\text{i}},\tau_{\text{f}},\Delta_{\text{i}},\Delta_{\text{f}})$ for each graph of the dataset. A noticeable observation is that the optimum parameters depend mostly on the hardness parameter and not on the size of the graph, see Fig. 1. The consequence is twofold: first, we can fit these data to generate a model of optimized schedules, as illustrated in Fig. 2. Second, such a model derived from small graphs can be extrapolated to larger graphs, see Sec. 3.2.
@@ 2.1 Piecewise linear protocol
Figure 1: Optimum parameters of the linear schedules for the 500 representative graphs. These values depend on $\mathcal{HP}$ and converge quickly to a constant for sufficiently hard graphs. This behavior is particularly strong for $\tau_{\text{f}}$ (top) and $\Delta_{\text{f}}$ (bottom) at the end of the adiabatic evolution. There is no statistical difference between the smallest graphs (down purple triangles) and the largest ones (up yellow-green triangles).
@@
Figure 2: Optimized piecewise linear schedules with 4 free variables. The Rabi amplitude (top) and the detuning (bottom) depend on the hardness of the graph and are illustrated here for $\mathcal{HP}\in\{0.5,1,2,4,8\}$ (thin lines, from green to maroon) and in the limit $\mathcal{HP}\to\infty$ (dark maroon thick line). The optimized values become in this limit $\tau_{\text{i}}\approx t_{\text{f}}/2$, $\tau_{\text{f}}\approx t_{\text{f}}/10$, $\Delta_{\text{i}}\approx-\Delta_{\text{UB}}/2$, and $\Delta_{\text{f}}=\Delta_{\text{UB}}$.
@@
Figure 3: Optimized piecewise linear schedules with 6 free variables. In this case, $\tau_{\text{i}}$ dislays a clear dependence on $\mathcal{HP}$. In the limit of infinite hardness, the optimized values are $\tau_{\text{i}}\approx 0.22\,t_{\text{f}}$ and $\tau_{\text{f}}\approx 0.12\,t_{\text{f}}$ for the Rabi amplitude (top) and $\Delta_{\text{i}}\approx-0.83\,\Delta_{\text{UB}}$, $\tau_{\text{m}}\approx 0.34\,t_{\text{f}}$, $\Delta_{\text{m}}\approx 0.36\,\Delta_{\text{UB}}$, and $\Delta_{\text{f}}=\Delta_{\text{UB}}$ for the detuning (bottom).
@@
The hardness-dependent linear schedules yield very good results compared to existing protocols, which motivates us to go one step further by considering two more degrees of freedom $(\tau_{\text{m}},\Delta_{\text{m}})$ for the detuning, see Fig. 3. The resulting optimized schedules are still easily implementable, yet they perform much better than all benchmark protocols, see Sec. 2.3.
A priori knowledge of the hardness parameter
The aim of the adiabatic evolution is to find the MIS of a graph, which in turn allows one to calculate the hardness of that very graph. In the case of small graphs, since all vertex configurations are enumerated effortlessly, we consider that $\mathcal{HP}$ is a known value that can be used to choose the best schedules from the fitted model. For large graphs, however, the hardness parameter cannot be considered to be known beforehand. In that case, one should either take the schedules in the limit $\mathcal{HP}\to\infty$, which yields quite good results for all graphs, or eventually apply a one-dimensional search on the hardness parameter, which is still much easier than performing a multi-dimensional optimization of all schedule parameters.
2.2 Counterdiabatic protocol

### v3

# Hardness-dependent quantum adiabatic schedules for the maximum-independent-set problem

@@ Abstract
We propose a numerical approach to design highly efficient adiabatic schedules for analog quantum computing, focusing on the maximum-independent-set problem and neutral atom platforms. On the basis of a representative dataset of small graphs, we present numerical evidence that the optimum schedules depend principally on the hardness of the problem rather than on its size. These schedules perform better than the benchmark protocols and admit a straightforward implementation in the hardware. This allows us to extrapolate the results to larger graphs and to successfully solve moderately hard problems using QuEra’s 256-qubit Aquila computer. We believe that extending our approach to hybrid algorithms could be the key to solving the hardest instances with the current technology, making yet another step toward real-world applications.
Maximum Independent Set; Hardness Parameter; Neutral Atoms; Rydberg Blockade; QuEra’s Aquila; Analog Quantum Computing; Adiabatic Schedules; Counterdiabatic Driving
I Introduction
@@
In recent years, quantum computers have emerged as a promising way to solve complex optimization tasks in science and industry, notably those based on combinatorics where an exhaustive search is not tractable [1, 2, 3]. Typically, the problem of interest is encoded in the ground state of a quantum Hamiltonian, which can be reached through adiabatic evolution starting from an easy-to-prepare initial state [4, 5, 6, 7]. Advanced protocols have to be developed to overcome not only the hardware limitations, such as qubit decoherence and control errors, but also the intrinsic complication of a vanishing gap in the eigenspectrum during the adiabatic evolution [8, 9, 10, 11].
@@
Here we focus on the maximum-independent-set (MIS) problem, where the goal is to find the largest set of vertices that do not share an edge in a graph. Such problems appear naturally in practical applications, for instance, network design [12] or financial analysis [13], and are known to be NP-hard in general [14]. Finding a MIS is still a difficult task in the particular case of unit disk graphs, in which any two vertices are connected by an edge if they are closer than a fixed distance [15, 16]. This family of graphs admits a direct mapping to arrays of Rydberg atoms, which explains the popularity of the MIS problem in analog quantum computing [17, 18, 19, 20, 21, 22, 23]. In that setting, optical tweezers enable arbitrary arrangements of atoms and controlled laser pulses manipulate the quantum states [24, 25]. The lasers then drive the evolution of the system according to schedules that are specific to the problem to be solved.
@@
The precise description of the schedules is of utmost importance since it conditions the success of the quantum algorithm. Many methods based on adiabatic criteria or on carefully chosen ansatzes can be applied in this case: quantum approximate optimization algorithm and variational quantum eigensolver [26, 27, 28], shortcut to adiabadicity and counterdiabatic (CD) protocol [29, 30, 31, 32], Bayesian and recursive optimization [33, 34], machine learning [35], and variational sampler [36], to name a few. Most of the aforementioned methods require multiple iterations to be run on a quantum device and to be updated by a classical optimizer and thus belong to the class of hybrid algorithms [37]. By their very nature, these algorithms cost both time and money as they make intense use of the quantum hardware, which motivates us to search for a more direct way of designing efficient schedules.
A detailed analysis of the classical dynamics of the simulated annealing for the MIS problem reveals that the graphs with many suboptimal solutions are likely to trap the algorithm in local minima, hence missing the global solution [19]. This observation leads to the definition of the hardness parameter $\mathbb{H}=D_{|\text{MIS}|-1}/(|\text{MIS}|\,D_{|\text{MIS}|})$, where $D_{k}$ is the degeneracy of the independent sets of size $k$. On the basis of the theory of Markov chains, $\mathbb{H}$ is shown to determine the performance of simulated annealing as it provides a lower bound on the inverse of its spectral gap and, by extension, on the hitting time for finding the solution. By analogy with the adiabatic theorem [6], this suggests that the hardness parameter may play a key role during the optimization of the schedules in the quantum setting, which is the starting point of our study.
In what follows, we show how a precomputation of optimized schedules for small graphs can lead to highly efficient protocols for graphs of larger size on the quantum device [38]. First, we review the MIS encoding in Rydberg atoms, emphasizing the importance of well-chosen boundary conditions. Then, in Sec. II we describe how to optimize the schedules on the basis of numerical simulations of small graphs. Finally, in Sec. III we present the results of cloud quantum computing demonstrations. Detailed information and calculations are provided in the appendices.
I.1 MIS encoding in Rydberg atom arrays
@@
The analog quantum dynamics of $N$ atoms is governed by the Rydberg Hamiltonian $H=H_{\text{drive}}+H_{\text{cost}}$, where the driving part and the cost function are [19]
@@
$\frac{H_{\text{cost}}(t)}{\hbar}$ $=-\Delta(t)\sum_{i=1}^{N}\hat{n}_{i}+\sum_{i<j}V_{ij}\hat{n}_{i}\hat{n}_{j}.$ (1b)
In these equations, $\Omega$ is the Rabi drive amplitude, $\phi$ its phase, $\Delta$ the detuning, and $\hat{n}_{i}=\ket{1}_{i}\bra{1}$ | detects the Rydberg excitation of the $i$th atom. The two-body potential is $V_{ij}=C_{6}/|\textbf{r}_{i}-\textbf{r}_{j}|^{6}$, with $C_{6}$ a constant for the van der Waals interaction between two Rydberg states at positions $\textbf{r}_{i}$ and $\textbf{r}_{j}$. Because of the large decay exponent, this interaction prevents the presence of any doubly excited state if two atoms are close together but has a negligible impact if they are further apart. This enables the encoding of MIS problems on unit disk graphs, where the so-called blockade radius of the Rydberg interaction corresponds to the maximum distance between any two connected vertices of the graph, as illustrated in Refs. [16, 39]; see Appendix A for the mathematical description of this encoding.
The boundary conditions at initial and final times $t_{\text{i}}$ and $t_{\text{f}}$ of the adiabatic quantum computation are usually stated as follows:
@@
In practice, we implement the first inequality as $\Delta_{\text{i}}\leq-\delta_{\Delta}$, where $\delta_{\Delta}$ is a small positive constant that incorporates the main sources of error in the quantum hardware, so that it is ensured that all atoms start in the easy-to-prepare ground state |$\ket{0}$. Regarding the MIS encoding, the second inequality turns out to be necessary but not sufficient, as the final detuning has to lie in a narrow interval that depends on the graph geometry, see Appendix A.2. For instance, the lower and upper bounds of the detuning for unit disk graphs on a square lattice are given by
$\Delta_{\text{LB}}=\frac{V_{0}}{12}\leq\Delta_{\text{f}}\leq\frac{V_{0}}{8}=\Delta_{\text{UB}},$ (3)
where $V_{0}=C_{6}/a^{6}$ is the interaction strength between adjacent nodes separated by distance $a$. Note that such bounds are not universal, but they are satisfied by the vast majority of graphs for a given lattice.
I.2 Quantum hardware
@@
In Aquila, the Rabi amplitude and the detuning must be defined as piecewise linear and not smooth functions, while the phase must be piecewise constant. However, the step size can be as small as 0.05 µs, which results in 80 intervals for the longest coherent evolution. The schedules are therefore well approximated in the quantum hardware, but one could slightly improve the results of the demonstrations by taking into account this hardware capability in the numerical simulations.
II Numerical optimization
We consider a fixed set of 500 representative graphs of size 8 to 17, which allows us to perform fast numerical simulations and to use efficient deterministic optimization algorithms, see Appendix C. This dataset is such that the correlation between the size of the graphs and their hardness parameter is minimized. Therefore, the results that depend principally on $\mathbb{H}$ can be extrapolated, to a certain extent, to graphs of any size.
@@ II Numerical optimization
where |$\ket{\psi(\Theta,t_{\text{f}})}$ is the final quantum state and |$\ket{\text{MIS}_{i}}$ is the basis vector that corresponds to the $i$th solution of the MIS problem. Note that most studies consider the approximation ratio instead (normalized mean energy of the final state), but in our opinion only the lowest energy level and not the full energy spectrum has practical utility because of the very definition of the hardness parameter.
Following the “best practices” listed in Ref. [25], we consider protocols for which the Rabi amplitude $\Omega(t)$ reaches $\Omega_{\max}$ during the evolution, so that the associated dynamic blockade radius $R_{\text{b}}$ becomes $(C_{6}/\Omega_{\max})^{1/6}\approx 8.4$. In order to generate a unit disk graph on the nearest and next-nearest neighbors of a square lattice, this radius should satisfy $\sqrt{2}a<R_{\text{b}}<2a$, and consequently $5.94.2<a<$ 5.9. We set $a=5$, which is close to the geometric mean of these bounds and which yields good results in preliminary calculations. We consider a fixed evolution time $t_{\text{f}}=1$ so that $P_{\text{MIS}}$ varies as much as possible depending on the various protocols; much shorter or longer times lead to either extremely bad or extremely good results independently of the precise shape of the schedules.
@@
II.1 Piecewise linear protocol
@@
Figure 1: Optimum parameters of the linear schedules for each graph of the dataset. These values depend on $\mathbb{H}$ and converge to a constant (dashed red lines) for sufficiently hard graphs. This behavior is particularly strong for $\tau_{\text{f}}$ (top) and $\Delta_{\text{f}}$ (bottom) at the end of the adiabatic evolution, with no statistical difference between the smallest graphs (purple triangles) and the largest ones (yellow-green triangles). We fit all points to extract a smooth and monotonic dependence on $\mathbb{H}$ (thick red lines).
The typical adiabatic protocol is based on piecewise linear schedules with two breakpoints, where $\Omega$ is slowly ramped up, then $\Delta$ is ramped from negative to positive, and finally $\Omega$ is ramped down; see Fig. 3.1b in Ref. [25] or Fig. 2c in Ref. [43]. Usually, these schedules are symmetric in the sense that the initial and final ramp times $\tau_{\text{i}}$ and $\tau_{\text{f}}$ are equal, and that the initial and final detunings $\Delta_{\text{i}}$ and $\Delta_{\text{f}}$ have opposite values. In our case, we are looking for schedules that are as general as possible, so we keep these parameters independent. Moreover, we do not impose the condition that the breakpoints of $\Omega$ coincide with those of $\Delta$.
After a few optimization steps, it appears that the flat portions of the detuning tend to vanish; we therefore consider a simple linear function for this schedule. Then we optimize the four remaining parameters $(\tau_{\text{i}},\tau_{\text{f}},\Delta_{\text{i}},\Delta_{\text{f}})$ for each graph of the dataset. A noticeable observation is that the optimum parameters depend mostly on the hardness parameter and not on the size of the graph. Consequently, we can fit these data to generate a model of optimized schedules, as illustrated in Figs. 1 and 2.
@@
Figure 2: Optimized piecewise linear schedules with four free variables. The Rabi amplitude (top) and the detuning (bottom) depend on the hardness of the graph and are illustrated here for $\mathcal{HP}\in\{0.5,1,2,4,8\}$ (thin lines, from green to maroon) and in the limit $\mathbb{H}\to\infty$ (thick dark maroon line). In this limit, the optimized values become $\tau_{\text{i}}\approx t_{\text{f}}/2$, $\tau_{\text{f}}\approx t_{\text{f}}/10$, $\Delta_{\text{i}}\approx-\Delta_{\text{UB}}/2$, and $\Delta_{\text{f}}=\Delta_{\text{UB}}$.
@@
Figure 3: Optimized piecewise linear schedules with six free variables. In this case, $\tau_{\text{i}}$ displays a clear dependence on $\mathbb{H}$. In the limit of infinite hardness, the optimized values are $\tau_{\text{i}}\approx 0.22t_{\text{f}}$ and $\tau_{\text{f}}\approx 0.12t_{\text{f}}$ for the Rabi amplitude (top) and $\Delta_{\text{i}}\approx-0.83\Delta_{\text{UB}}$, $\tau_{\text{m}}\approx 0.34t_{\text{f}}$, $\Delta_{\text{m}}\approx 0.36\Delta_{\text{UB}}$, and $\Delta_{\text{f}}=\Delta_{\text{UB}}$ for the detuning (bottom).
@@
The hardness-dependent linear schedules yield very good results compared with existing protocols, which motivates us to go one step further by considering two more degrees of freedom $(\tau_{\text{m}},\Delta_{\text{m}})$ for the detuning; see Fig. 3. The resulting optimized schedules are still easily implementable, yet they perform much better than all benchmark protocols; see Sec. II.3.
Let us now provide some hints about the performance of the hardness-dependent linear schedules. The adiabatic condition for quantum computing is traditionally given by [6]
$t_{\text{f}}\gg\max_{s\in[0,1]}\frac{|\bra{\varepsilon_{1}(s)}\partial_{s}H(s)\ket{\varepsilon_{0}(s)}|}{|\varepsilon_{1}(s)-\varepsilon_{0}(s)|^{2}}=t_{\text{ad}},$ (5)
where $s=t/t_{\text{f}}$ is the dimensionless time, |$\ket{\varepsilon_{0}(s)}$ is the instantaneous ground state of $H(s)$ with energy $\varepsilon_{0}(s)$, and |$\ket{\varepsilon_{1}(s)}$ is any of the first excited states; for simplicity, we consider only the nondegenerate MIS instances here. In particular, Eq. (5) implies that the minimum energy gap $\Gamma=\min_{s}\varepsilon_{1}(s)-\varepsilon_{0}(s)$ should be large to maintain adiabaticity. Unsurprisingly, this gap gets smaller as the hardness parameter increases: from the numerical simulations of the linear schedules with six parameters, we find $\Gamma\propto\exp(-\alpha\mathbb{H}^{\beta})$ with $\alpha\approx 0.76$ and $\beta\approx 0.36$. In this respect, it seems counterintuitive that the Rabi amplitude should be ramped up and down more rapidly for harder graphs; see Fig. 3. Faster variations are likely to introduce stronger diabatic transitions which, in this case, would not be compensated by a larger gap to satisfy the adiabatic condition. However, we show in Fig. 4 that the optimized schedules do minimize the timescale $t_{\text{ad}}$ on average, which confirms that adiabaticity lies at the heart of the protocol’s performance.
Figure 4: Adiabatic criteria for the $\mathbb{H}$-linear protocol with six parameters. Left: Minimum energy gap for “easy” ($\mathbb{H}\leq 4$) and “hard” ($\mathbb{H}>4$) graphs, based on the schedules illustrated in Fig. 3 with $\mathbb{H}=1$ (light green) and $\mathbb{H}\to\infty$ (maroon). The gap by itself does not explain the hardness dependence of the optimum schedules, as shorter ramp times of the detuning (maroon boxes) are slightly beneficial for all graphs. Right: Corresponding adiabatic timescales. In this case, $t_{\text{ad}}$ is indeed minimized if one applies the schedules according to the hardness of the graphs (leftmost and rightmost boxes).
II.2 Counterdiabatic protocol